OpenStudy (anonymous):

are these numbers rational: sqt16 sqt -15 pi 1.765 -1.5 sqt-6 0 e

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

1 = rational. 2 = irrational.

OpenStudy (saifoo.khan):

3 = ir 4 = ir 5 = ir 6 = ir

OpenStudy (saifoo.khan):

idk about 7th & 8th.

OpenStudy (anonymous):

4 is rational

OpenStudy (anonymous):

2 is irrational?

OpenStudy (saifoo.khan):

\[4th = \frac{353}{200}\]

OpenStudy (anonymous):

These are rational 0 sqrt16 1.765 -1.5

OpenStudy (saifoo.khan):

Yes, 2nd is. coz it contains imaginory(sp) terms

OpenStudy (saifoo.khan):

-1.5 = -3/2 irrational.

OpenStudy (anonymous):

-3/2 is a quotient of two integers , so rational

jimthompson5910 (jim_thompson5910):

-1.5 = -3/2 is rational (there's a fraction of two integers)

OpenStudy (saifoo.khan):

but it's in fraction.

OpenStudy (saifoo.khan):

Might be.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Rational_number

OpenStudy (saifoo.khan):

on it.